# Compositional amortized inference for large-scale hierarchical Bayesian models

### v1

@@ 3.3 Scalable compositional score matching
We propose using different noise schedules for training and inference. During inference, spending less time in the high-noise region improves stability and allows for larger step sizes, which is particularly important in the large-$J$ regime. In the case of the cosine schedule $\lambda(t)=-2\cdot\log(\tan(\pi t/2))+2s$ proposed by Nichol and Dhariwal [2021], this can be easily achieved by increasing the shift parameter $s$, which effectively compresses the high-noise segment of the schedule. As discussed by Karras et al. [2022] and Kingma and Gao [2023], the combination of the noise schedule and the weighting function in the denoising score matching objective plays a role akin to importance sampling.
Figure 2: Evaluation of the error-damping estimator for the toy example. Different evaluation metrics are shown for different data set sizes and damping factors $d_{1}$ or cosine sifts $s$. The mini-batch size was set to 10% of the data set size and for each step 10 runs were performed. The median and median absolute deviation is reported besides for those runs where none converged.
@@ 4 Experiments
  • Hierarchical time series model: A grid of AR(1) processes with shared global and local parameters, used to evaluate joint inference performance against MCMC.
@@
4.1 Experiment 1: scaling and stabilizing CSM with error-damping mini-batch estimation
This first experiment serves both as a sanity check and as a demonstration of the stabilizing effects of our error-damping estimator, highlighting the accuracy and scalability of compositional score matching in a controlled setting. We consider a Gaussian model of dimension $D{=}10$ with conditionally independent groups and a global latent variable (see Appendix A.5). Since the posterior is analytically tractable, it enables precise measurement of accuracy and convergence. We scale the number of observations up to 100,000 to test the effect of dataset size on error accumulation of the individual scores. Below, we summarize our results and provide practical recommendations.
@@ 4.1 Experiment 1: scaling and stabilizing CSM with error-damping mini-batch estimation
We find that the optimal damping factor $d_{1}$ depends on the number of composed groups: larger datasets require smaller damping factors for convergence (Figure 2). However, overly small factors can prevent posterior contraction, worsen calibration, and even hinder convergence. With an initial factor of $0.003$, we successfully compose 100,000 scores. At this scale, the analytical posterior becomes nearly a point estimate, so even slight deviations in our estimate can significantly increase the KL-divergence, but the RMSE remain negligible. The damping factor is a tunable hyperparameter, and a value on the order of $1/n$ often serves as a good starting point.
@@
Our mini-batch estimator reduces computational cost per sampling step but does not resolve instability due to score error accumulation, which prevents convergence beyond 1000 groups (see Appendix Figure 7). Using smaller batches instead of the full dataset lowers both the KL-divergence and posterior calibration error, albeit with a slight increase in RMSE. We attribute this to a smoothing effect on accumulated score errors. In practice, we recommend using mini-batches of about 10% of the data to balance robustness and computational demands.
Table 2: Benchmarking against NUTS (gold-standard MCMC) for the hierarchical AR(1) model.
@@
Adjusting the noise schedule improves stability and mitigates error accumulation (Appendix Figure 7). A large shift of $s{=}10$ enables scaling to 100,000 groups and improves KL-divergence, RMSE, and calibration error. As expected, both KL-divergence and calibration degrade with larger datasets due to increased error accumulation, but the shifted schedule helps to mitigate this effect. Moreover, linear or EDM sampling schedules appear suboptimal for compositional score matching, failing to converge even on smaller datasets (see Appendix Figure 8).
@@
Computing joint scores based on groups of 100 groups allows inference on up to 100,000 groups (Appendix Figure 7). However, increasing the number of conditioning groups does not necessarily lead to better posterior contraction or lower RMSE. Notably, the number of conditions has to be chosen before training, and conditioning on more groups requires additional simulations, since each training sample incorporates multiple groups. Further, we find that the latter demand better network capacity, otherwise, RMSE deteriorates. In practice, using a small number of conditions can yield performance gains without incurring major training cost.
In summary, our experiment with the analytically tractable Gaussian toy example demonstrates that the error-damping mini-batch estimator affords scalable compositional inference for up to 100,000 groups. While mini-batching alone is insufficient to ensure convergence, combining it with damping and noise schedule shifting reduces score accumulation errors and computational cost.
@@ 4.2 Experiment 2: scaling hierarchical Bayesian inference
Our second experiment evaluates whether compositional score matching can accurately infer both global and local parameters for a non-trivial hierarchical model. We simulate a grid of local AR(1) processes with a shared global drift and local variation parameters (see Appendix A.6). We increase the grid size from $4{\times}4$ to $128{\times}128$ to test the scalability of the method, resulting in up to 16,384 local latent vectors.
Figure 3: Assessing inference for high-resolution grids ($128{\times}128$). A Global parameter recovery across datasets, showing the posterior median and median absolute deviation. B Posterior calibration plot for the global parameters, using the graphical diagnostic method from Säilynoja et al. [2022].
Figure 4: Evaluation of the error-damping estimator for the hierarchical AR(1) model. Different evaluation metrics are shown for different data set sizes and damping factors $d_{1}$.
@@
For this grid of AR(1) processes direct comparison to NUTS [as implemented in Stan Carpenter et al., 2017] is possible, which is widely regarded as a gold standard for Bayesian inference and provides the most reliable benchmark for evaluating how well our method captures the correct shrinkage in the local parameters. Thus, we compare our approach with Stan on grids up to a resolution of $32{\times}32$. At this resolution, posterior sampling with Stan for 100 datasets takes a similar amount of time as training one score-based model and performing amortized inference across the same datasets.
@@
Our experiments with the hierarchical model on the damping factor support the earlier findings: tuning the damping function is essential to balance posterior contraction and estimation error (Figure 4), however a too large cosine shift might hinder calibration. Moreover, we find that neither the damping factor nor the cosine shift alone is sufficient to ensure convergence on high-resolution grids (e.g., $128{\times}128$), but their combination stabilizes inference (Figure 3-4). However, for these large-scale settings, achieving well-calibrated posteriors often comes at the cost of reduced accuracy in parameter recovery. This difficulty arises due to the strong contraction of the posterior in hierarchical models. As a result, calibration becomes especially challenging in the high-resolution regime.
@@
We observe that our method yields results comparable to NUTS at both the global and local parameter levels (Table 2; showing mean and standard deviation over parameters), with a bit higher RMSE on the local level. Crucially, our method scales effortlessly to significantly larger grid sizes, such as $128{\times}128$ (Figure 3). In contrast, Stan becomes computationally infeasible at this scale: inference for 100 datasets on such grids would require approximately 16 hours, whereas our approach completes inference within a few minutes using the learned score model.
In summary, our experiment with the grid of AR(1) processes shows that compositional score matching, when combined with damping and noise schedule shifting, enables accurate and scalable inference in hierarchical models with thousands of groups. Our method performs comparably to MCMC on small grids and remains stable on large grids where MCMC becomes computationally infeasible. While strong posterior contraction introduces challenges for calibration at scale, the proposed approach offers a practical and efficient approach for amortized hierarchical inference.
Figure 5: Inference for fluorescence lifetime imaging. A Mean intensity across time for each pixel, representing the fluorescence data. B Time series data and fitted posterior median for representative pixels. C Spatial map of the fitted local posteriors (medians) across the full image and mean lifetime.
@@ 4.3 Experiment 3: application to fluorescence lifetime imaging (FLI)
Fluorescence lifetime imaging (FLI) is an important tool in pre-clinical cancer imaging, particularly for in vivo drug-target analysis [Verma et al., 2025]. However, FLI remains challenging due to its reliance on time-resolved acquisition at sub-nanosecond resolution, computationally intensive pixel-wise curve fitting, and noisy decay profiles caused by low quantum yield dyes, resulting in high uncertainty in lifetime estimates [Yuan et al., 2024, Trinh and Esposito, 2021]. A Bayesian approach has been explored in prior work [Wang et al., 2019, Rowley et al., 2016], but to our knowledge, this is the first application of a fully hierarchical Bayesian model to FLI data.
We analyze time-resolved fluorescence decay data from FLI (Figure 5A). Therefore, each pixel in the measured $512{\times}512$ images is modeled using a bi-exponential decay with local decay rates and mixture weights (Figure 5B), resulting in a hierarchical inference problem with over 250,000 groups. Unlike amortized methods that require full-image simulations to generalize across spatial structures, our approach trains on pixel-wise simulations, requiring only the equivalent of 350 images.
To assess the performance of the standard non-hierarchical approach and our proposed method, we first consider synthetic FLI data. We found that per-pixel maximum likelihood estimation (MLE) fails to recover the ground truth due to photon-limited noise. In contrast, our hierarchical approach accurately captures both global and local structures (Appendix 4.3). Nevertheless, estimating global variances remains challenging under very high noise conditions.
Encouraged by these results, we applied the method to real FLI data (details in Appendix 4.3). Using the trained score-based hierarchical model, we efficiently and reliably fit over 750,000 local parameters (Figure 5C). The inferred mean lifetime closely matches that of a standard MLE fit (Appendix Figure 13), which is commonly recovered by conventional methods [Torrado et al., 2024]. Moreover, the model achieves excellent fits across the full image, as illustrated by representative decay curves and their posterior median reconstructions (Figure 5B).
In summary, this experiment demonstrates the scalability and practical utility of our approach for real-world data, enabling stable posterior estimates where traditional methods struggle with noise or high dimensionality.
@@ 5 Conclusion
HM are of outstanding importance, but their inference remains challenging. Here, we demonstrated that compositional score matching (CSM) provides a scalable and flexible framework for estimating non-trivial hierarchical Bayesian models. Moreover, we introduced an error-damping mini-batch estimator that resolves the inherent instability of CSM up to hundreds of thousands of data points. As a notable limitation, we observed that posterior calibration becomes difficult at scale, particularly under extreme contraction. While our dampening strategy reduces instability, achieving good calibration across all levels is an open problem. Future work could further explore temporal aggregation [Gloeckler et al., 2024b], systematically test the trade-off introduced by different dampening schedules, refine mini-batch selection using informativeness criteria [Peng et al., 2019], and generalize to more than two levels.

### v3

@@ 3.2 Stabilizing and scaling compositional score matching
Finally, we propose to use different noise schedules for training and inference. During inference, spending less time in the high-noise regime of the reverse process improves stability and allows for larger step sizes, which is particularly important in the large-$J$ setting. In the case of the cosine schedule $\lambda(t)=-2\cdot\log(\tan(\pi t/2))+2s$ proposed by Nichol and Dhariwal [2021], this can be easily achieved by increasing the shift parameter $s$, which effectively compresses the high-noise portion of the schedule. As discussed by Karras et al. [2022] and Kingma and Gao [2023], the combination of the noise schedule and the weighting function in the denoising score matching objective plays a beneficial role akin to importance sampling.
3.3 Compositional score matching for hierarchical modeling
To employ our stable compositional formulation for simulation-efficient hierarchical Bayesian modeling, we represent the posterior at each hierarchical level with its own score estimator. The outputs of these score estimators are then connected via the inverse factorization (see Figure 1). This design is similar to the frameworks introduced by Habermann et al. [2024] and Heinrich et al. [2023] but avoids the need for hierarchical embeddings and exhaustive simulation. At higher levels of the hierarchy, we use our stable compositional formulation (Eq. 9), enabling training of the global score model on single groups. For example, in a two level model, we have
$s_{\boldsymbol{\psi}}^{\text{local}}(\boldsymbol{\theta}_{t,j},\boldsymbol{\eta},\mathbf{Y}_{j},\lambda_{t})\approx\nabla_{\boldsymbol{\theta}_{t}}\log p_{t}(\boldsymbol{\theta}_{t,j}\mid\boldsymbol{\eta},\mathbf{Y}_{j}),\,s_{\boldsymbol{\psi}}^{\text{global}}(\boldsymbol{\eta}_{t},\mathbf{Y}_{j},\lambda_{t})\approx\nabla_{\boldsymbol{\eta}_{t}}\log p_{t}(\boldsymbol{\eta}_{t}\mid\mathbf{Y}_{j}).$ (10)
For each group, we may learn a shared summary representation $\mathbf{h}_{j}=h(\mathbf{Y}_{j})$ via a summary network $h$. Either the raw data $\mathbf{Y}_{j}$ or its summary $\mathbf{h}_{j}$ is then used as input to both the global and local score-based models. The design of the summary $h$ should be adapted to the specific data modality (e.g., recurrent networks or transformers for time series, etc.). When conditioning on multiple groups, we encode exchangeability via a second summary network [e.g., a DeepSet, Zaheer et al., 2017], which aggregates the individual summaries into a permutation-invariant global summary.
The global and local score networks can be trained jointly via denoising score matching objectives,
$\min_{\boldsymbol{\psi}}\mathbb{E}_{p(\boldsymbol{\theta},\boldsymbol{\eta},\mathbf{Y})}\mathbb{E}_{t\sim\mathcal{U}(0,1)}w_{t}\left[\|\boldsymbol{\epsilon}+s_{\boldsymbol{\psi}}^{\text{local}}(\boldsymbol{\theta}_{t},\boldsymbol{\eta},\mathbf{Y},\lambda_{t})\sigma_{t}\|_{2}^{2}+\|\boldsymbol{\epsilon}+s_{\boldsymbol{\psi}}^{\text{global}}(\boldsymbol{\eta}_{t},\mathbf{Y},\lambda_{t})\sigma_{t}\|_{2}^{2}\,\right],$ (11)
with $\boldsymbol{\eta}_{t}=\alpha_{t}\boldsymbol{\eta}+\sigma_{t}\boldsymbol{\epsilon}$ and $\boldsymbol{\theta}_{t}=\alpha_{t}\boldsymbol{\theta}+\sigma_{t}\boldsymbol{\epsilon}$, where $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. Having trained the score models, we can sample from the joint posterior via ancestral sampling:
$\boldsymbol{\eta}$ $\sim q_{\boldsymbol{\psi}}^{\text{global}}(\boldsymbol{\eta}\mid\{\mathbf{Y}_{j}\}_{j=1}^{J}),\qquad\boldsymbol{\theta}_{j}\sim q_{\boldsymbol{\psi}}^{\text{local}}(\boldsymbol{\theta}\mid\boldsymbol{\eta},\mathbf{Y}_{j}),$ (12)
where we use the compositional score (Eq. 10) to sample the global parameters and then sample the local parameters conditioned on the global sample using standard score-based diffusion.
Figure 2: Evaluation of the error-damping estimator for the Gaussian toy example. Different evaluation metrics are shown for different data set sizes and damping factors $d_{1}$ or cosine shifts $s$. The mini-batch size was set to 10% of the data set size and for each step 10 runs were performed. The median and median absolute deviation is reported, besides for those runs where none converged.
@@ 4 Experiments
  • Hierarchical time series model: A grid of AR(1) processes with shared global and local parameters, used to evaluate hierarchical estimation against gold-standard MCMC.
@@
4.1 Experiment 1: Scaling and stabilizing CSM with error-damping estimation
Setup and baseline
This first experiment serves both as a sanity check and as a demonstration of the stabilizing effects of our error-damping estimator, highlighting the accuracy and scalability of compositional score matching in a controlled setting. We consider a Gaussian model of dimension D=10D{=}10 with conditionally independent groups and a global latent variable (see Appendix A.5). Since the posterior is analytically tractable, it enables exact measurement of accuracy and convergence. We scale the number of observations up to 100,000 to test the effect of dataset size on error accumulation of the individual scores. Below, we summarize our results and provide practical recommendations.
@@ 4.1 Experiment 1: Scaling and stabilizing CSM with error-damping estimation
We find that the optimal damping factor $d_{1}$ depends on the number of composed groups: larger datasets require smaller damping factors for convergence (Figure 2). However, overly small factors can prevent posterior contraction, worsen calibration, and even hinder convergence. With an initial factor of $0.1$, we can successfully compose 100,000 scores. At this scale, the analytical posterior becomes nearly a point estimate, so even slight deviations in our estimate can significantly increase the KL-divergence, but the RMSE remain negligible. The damping factor is a tunable hyperparameter, and a value on the order of $1/\sqrt{n}$ often serves as a good starting point.
@@
Our mini-batch estimator reduces computational cost per sampling step but does not resolve instability due to score error accumulation, which prevents convergence beyond 1000 data points (see Appendix Figure 6). Using smaller batches instead of the full dataset lowers both the KL-divergence and posterior calibration error, albeit with a slight increase in RMSE. We attribute this to a smoothing effect on accumulated score errors. In practice, we recommend using mini-batches of about 10% of the data to balance accuracy and computational demands.
Table 2: Benchmarking against NUTS (gold-standard MCMC) for the hierarchical AR(1) model. We show mean and standard deviation over parameters.
@@
Adjusting the noise schedule improves stability and mitigates error accumulation (Appendix Figure 6). A large shift of s=10s{=}10 enables scaling to 100,000 groups and improves KL-divergence, RMSE, and calibration error. As expected, both KL-divergence and calibration degrade with larger datasets due to increased error accumulation, but the shifted schedule helps to mitigate this effect. Moreover, linear or EDM sampling schedules appear suboptimal for compositional score matching, failing to converge even on smaller datasets (see Appendix Figure 7).
@@
Scaling to 100,000 groups becomes also feasible by conditioning the diffusion model on subsets of 100 groups (Appendix Figure 6). However, increasing the number of conditioning groups does not necessarily lead to better posterior contraction or lower RMSE. Notably, the number of conditions has to be chosen before training, and conditioning on more groups requires additional simulations, since each training sample incorporates multiple groups. The choice of the number of groups per subset introduces a trade-off between scalability, accuracy, and the required expressivity of the summary network. While larger subsets can reduce the variance in the compositional score estimation, they require more expressive networks to compose group-level information into accurate score estimates. In practice, using a small number of conditions can yield performance gains without incurring major training costs.
In summary, our experiments with the analytically tractable Gaussian toy example demonstrates that the error-damping mini-batch estimator affords scalable compositional inference for up to 100,000 units of information. While mini-batching alone is insufficient to ensure convergence, combining it with damping and noise schedule shifting reduces score accumulation errors and computational cost.
@@ 4.2 Experiment 2: Scaling hierarchical Bayesian inference
Setup and baseline
@@
Our second experiment evaluates whether our approach can accurately infer the joint posterior for a non-trivial hierarchical Bayesian model. We simulate a grid of local AR(1) processes with a shared global drift and local variation parameters (see Appendix A.6). We increase the grid size up to $128{\times}128$ to test the scalability of the method, resulting in up to 16,384 local parameter vectors. For this grid of AR(1) processes, direct comparison to NUTS [as implemented in Stan, Carpenter et al., 2017] is possible, which is widely regarded as a gold standard for Bayesian inference and provides the most reliable benchmark for evaluating how well our method captures the correct shrinkage in the local parameters. Importantly, our total simulation budget amounts to 610 grids of size $128\times 128$, rendering previous amortized methods that train on full grids completely infeasible with this low number of training samples.
Figure 3: Assessing inference for high-resolution grids ($128{\times}128$). A Global parameter recovery across 100 datasets, showing the posterior median and median absolute deviation. B Posterior calibration plot for the global parameters, using simulation-based calibration [Säilynoja et al., 2022].
Results
@@
Our results support the earlier findings regarding the role of the damping factor: tuning the damping function is essential to balance posterior contraction and estimation error (Appendix Figure 8), however a too large cosine shift might hinder calibration. Moreover, we find that neither the damping factor nor the cosine shift alone is sufficient to ensure convergence on high-resolution grids (e.g., $128{\times}128$), but their combination stabilizes inference (Figure 3). However, for these large-scale settings, achieving well-calibrated posteriors often comes at the cost of reduced accuracy in parameter recovery. This difficulty arises due to the strong contraction of the global posterior and compounding errors while solving the reverse SDE. As a result, calibration becomes challenging in the high-resolution regime.
@@
In terms of precision, we observed that our method yields results comparable to NUTS at both the global and local parameter levels (Table 2), with a slightly higher local RMSE. Crucially, our method scales effortlessly to significantly larger grid sizes, such as $128{\times}128$ (Figure 3). In contrast, NUTS requires approximately 9 hours on a high-performance cluster with 64 CPU cores, whereas our likelihood-free approach completes inference within a few minutes on a single GPU. Moreover, already at a resolution of $32{\times}32$, posterior sampling with NUTS for 100 datasets takes a similar amount of time as training one score-based model and performing amortized inference.
Inference-time hyperparameter optimization
Importantly, because inference with our method is amortized, we can perform grid-based or even Bayesian hyperparameter optimization. We tuned the damping factor and noise shift by selecting the best configuration based on the sum of the RMSE and calibration error. To generalize beyond our proposed decay damping, we introduce a flexible decay function: $d(t)=d_{0}+(d_{1}-d_{0})\cdot\left(1-(1-t^{\alpha})^{\beta}\right),$ which adds two hyperparameters ($\alpha$ and $\beta$) that enable smooth interpolation between linear, exponential-like, and cosine-like behaviors. We perform Bayesian optimization over $\alpha,\beta\in[0.3,2]$, $d_{1}\in[10^{-5},10^{-1}]$, and $d_{0}\in[10^{-3},1]$. This increases the runtime on a $32\times 32$ grid from 3 to 7 minutes, primarily due to a reduction in early failures during sampling. The best configuration yielded $d_{1}=0.005$, $d_{0}=0.94$, $s=3.53$, $\alpha=0.39$, and $\beta=1.97$, suggesting that the learned schedule strongly favors a sharp, exponential-like decay (Figure 3).
In summary, our experiment with the hierarchical AR(1) model revealed that compositional score matching, when combined with damping and noise schedule shifting, enables accurate and scalable inference in hierarchical models with thousands of groups. Even though NUTS is competitive on small grids, its cost and requirement for a tractable likelihood can make it impractical for estimating complex models from large data sets, whereas our compositional approach remains viable.
@@ 4.3 Experiment 3: Application to fluorescence lifetime imaging (FLI)
Practical relevance
Our final experiment demonstrates the practical utility of our approach for real-world data, enabling scalable posterior estimation in fluorescence lifetime imaging (FLI), where existing methods struggle with noise and high dimensionality. FLI is an important tool in pre-clinical cancer imaging, particularly for in vivo drug-target analysis [Verma et al., 2025]. However, FLI remains challenging because it requires sub-nanosecond time-resolved acquisition, computationally heavy pixel-wise curve fitting, and must deal with noisy decay profiles from low-quantum-yield dyes, leading to high uncertainty [Yuan et al., 2024, Trinh and Esposito, 2021]. Bayesian approaches have been explored in prior work [Wang et al., 2019, Rowley et al., 2016], but, to the best of our knowledge, we present the first application of a fully hierarchical Bayesian model to FLI data.
Figure 4: Inference for fluorescence lifetime imaging. A Mean intensity across time for each pixel, representing the fluorescence data. B Time series data and fitted posterior median for representative pixels. C Spatial map of the fitted local posteriors (medians) per pixel. D Spatial map of the coefficient of determination for each pixel, comparing our results with a popular baseline (MLE).
Setup and baseline
We analyze time-resolved fluorescence decay data (Figure 4A-B), where each pixel in a measured series of $512{\times}512$ images is modeled using a bi-exponential decay with local decay rates $\tau_{1}^{L}$ and $\tau_{2}^{L}$ and mixture weights $A^{L}$. Each local parameter has a global mean and a global standard deviation, resulting in a hierarchical inference problem with over 250,000 groups (see Appendix A.7). Unlike amortized methods that train on full-image simulations to generalize across spatial structures, our approach trains on single pixels, requiring only the equivalent of 350 full images for training. We compare our approach with the field’s gold standard method based on maximum likelihood estimation (MLE).
Results
To assess the performance of the baseline non-hierarchical approach and our proposed method, we first consider 100 held-out synthetic images. We found that per-pixel MLE fails to recover the ground truth due to photon-limited noise. In contrast, our hierarchical approach accurately captures both global and local structures (Appendix Figure 9-10). Nevertheless, estimating global variances remains challenging under very high noise conditions. Finally, we applied our method to real FLI data (Appendix A.7). Using the trained score-based hierarchical model, we fitted over $750,000$ local parameters efficiently (Figure 4C). Qualitatively, the inferred mean lifetime closely matches a standard MLE fit (Appendix Figure 11). Our approach achieves excellent image-wide fits, with mean $R^{2}=0.961$ (s.d., 0.017) for posterior predictive medians, versus $0.871$ (s.d., 0.110) for MLE (Figure 4D), as illustrated in Figure 4B. Across pixels, the mean posterior predictive $p$-value is 0.20 (s.d., 0.337), indicating slight underdispersion; masking the final third of the decay tail increases the mean $p$-value to 0.40 (s.d., 0.38), confirming that our model captures the core dynamics.
@@ 5 Conclusion
Hierarchical Bayesian models (HBMs) are of utmost importance in statistics, but their estimation remains challenging. Here, we demonstrated that compositional score matching (CSM) provides a scalable and flexible framework for estimating large HBMs. Moreover, we introduced an error-damping mini-batch estimator that resolves the inherent instability of CSM up to hundreds of thousands of data points. As a notable limitation, we observed that posterior calibration becomes difficult at scale, particularly under extreme contraction. Future work could further explore temporal aggregation [Gloeckler et al., 2024b], systematically test the trade-off introduced by different damping schedules, refine mini-batch selection using informativeness criteria [Peng et al., 2019], and generalize our experiments to more than two levels.